I HAVE A COUPLE QUESTIONS HERE:
#1 Find the slope, x, and y- intercepts for x = -3
The options to choose from are:
A) No Slope; (-3, 0)
B) No Slope
C) m = 0; (-3, 0)
#2 Find the slope, x, and y- intercepts for -1/3x - 2/3y = 5/3
Options are:
A) m = -2; (-5, 0), (0, 5/2)
B) m = 1/2; (0, 5/2), (0, -5)
C) m = -1/2; (-5, 0), (0, 5/2)
#3 Find the slope, x, and y- intercepts of (A/D)x + (B/D)y = C/D
Options are:
A) m = A/B; (C/A, 0), (0, C/B)
B) m = -A/B; (0, C/A), (C/B, 0)
C) m = -A/B; (C/A, 0), (0, C/B)

Answer by Alan3354(69443)

#1 Find the slope, x, and y- intercepts for x = -3
It's a line parallel to the y-axis thru the point (-3,0)
The slope is undefined or infinite (not "no slope")
x-int = (-3,0)
No y-intercept
----------------
A) No Slope; (-3, 0)
B) No Slope
C) m = 0; (-3, 0)
#2 Find the slope, x, and y- intercepts for -1/3x - 2/3y = 5/3
-1/3x - 2/3y = 5/3
-x - 2y = 5
Solve for y
y = (-1/2)x - 5/2
m = -1/2
x-int = (-5,0)
y-int = (0,-5/2)
--------------
A) m = -2; (-5, 0), (0, 5/2)
B) m = 1/2; (0, 5/2), (0, -5)
C) m = -1/2; (-5, 0), (0, 5/2)
#3 Find the slope, x, and y- intercepts of (A/D)x + (B/D)y = C/D
Ax + By = C
y = (-A/B)x + C/B
m = -A/B
x-int = (C/A,0)
y-int = (0,C/B)
A) m = A/B; (C/A, 0), (0, C/B)
B) m = -A/B; (0, C/A), (C/B, 0)
C) m = -A/B; (C/A, 0), (0, C/B)